OpenStudy (anonymous):

who wants medal!!!! =) An artist wants to make alabaster pyramids using a block of alabaster with a volume of 576 cubic inches. She plans to make each pyramid with a square base area of 3 square inches and a height of 4 inches. At most, how many pyramids can the artist make from the block of alabaster? 136 144 154 166 172

OpenStudy (anonymous):

Volume of Pyramid = area of base * height * 1/3 Volume of Pyramid = 3 * 4 * 1/3 = 4 ins^3 The artist can make 576/4 = 144 pyramids out of the block of alabaster.

OpenStudy (igreen):

@bajji You got that from here: http://www.algebra.com/cgi-bin/Surface-area.faq.question.795311.html

OpenStudy (anonymous):

It does not matter where you get the answer from it only matters wether you know how to go about the problem and if there is any doubt I can clarify it! :) Saves time right? :)! @iGreen

OpenStudy (igreen):

@bajji \(\hspace{80pt}\bbox[5pt, black,border:5px solid #00FF00]{\bf\huge \color{#00FF00}{\href{javascript:confirm('Your OpenStudy user has been issued a warning by iGreen. Please site your sources. refer to our code of conduct at openstudy.com/code-of-conduct for more information.')}{Click~Here}}}\)

OpenStudy (igreen):

Yes, but you should site your sources :)

OpenStudy (anonymous):

Funny because I clicked that, oops....

OpenStudy (igreen):

You just plagiarized. @bajji

OpenStudy (anonymous):

@iGreen i REALLY didnt know that! :O :) Okay... will site my sources from next time! :) Thank You! @iGreen

OpenStudy (igreen):

No problem :)

OpenStudy (anonymous):

thanks!!! =)